### v1

# Field of View Enhanced Signal Dependent Binauralization with Mixture of Experts Framework for Continuous Source Motion

Abstract
We propose a novel mixture of experts framework for field-of-view enhancement in binaural signal matching. Our approach enables dynamic spatial audio rendering that adapts to continuous talker motion, allowing users to emphasize or suppress sounds from selected directions while preserving natural binaural cues. Unlike traditional methods that rely on explicit direction-of-arrival estimation or operate in the Ambisonics domain, our signal-dependent framework combines multiple binaural filters in an online manner using implicit localization. This allows for real-time tracking and enhancement of moving sound sources, supporting applications such as speech focus, noise reduction, and world-locked audio in augmented and virtual reality. The method is agnostic to array geometry offering a flexible solution for spatial audio capture and personalized playback in next-generation consumer audio devices.
Index Terms—  Spatial Audio, Beamforming, Microphone Arrays, Binaural Rendering, Mixture of Experts
1 Introduction
Consumer audio capture devices are increasingly designed as wearable technologies. Among these, headworn microphone arrays have gained significant attention for capturing sound fields and enabling binaural rendering. A key use case arises when the user wishes to re-experience the recording in a way that matches how it originally sounded. This places importance on downstream processing methods that preserve the auditory cues present at the time of capture [1]. The process of filtering and summing the microphone array signals to reproduce the binaural cues at the left and right ears of the user has been referred to as end-to-end magnitude least squares (eMagLS) [2] or binaural signal matching (BSM) [3]. An alternative method to render binaural signals relies on direction-of-arrival (DOA) estimation and beamforming to extract direct signal components and renders direct and reverberant sound field components separately [4]. It is desirable to give the user additional control over properties of the rendered audio. Such control may take the form of speech enhancement, noise reduction, support for world-locked audio where playback adapts to user movement, or directional enhancement where sounds from selected directions or from the user’s gaze are emphasized [5, 6, 7]. This paper develops a signal-dependent framework for the latter task, which we refer to as field-of-view enhancement (FoVE) [8]. Moreover, we propose a mixture of experts algorithm that is able to combine the estimates of numerous signal-dependent binaural signal matching filters in an online manner with implicit localization instead of relying on traditional direction of arrival estimators. This allows the model to track a continuous talker without assumptions of stationarity. Figure 1 depicts the system diagram for the proposed method.
Fig. 1: The proposed binauralization method. Each directional output of the beamspace transform is rendered using signal-dependent binaural signal matching. Each candidate direction is deemed an ”expert” and the methods seeks to adaptively blend the estimates from different experts.
The objective of FoVE is to maintain the spatial structure of the recorded sound field while allowing the user to manipulate the directional response of the binaural rendering filter. In practice, this means enabling the user to increase or decrease the relative energy from chosen directions in a perceptually consistent manner. This produces emphasis of sources in the selected region and suppression of sources outside it, while preserving binaural cues and spatial consistency across directions. Importantly, this is not restricted to speech sources but applies to more general sound fields. The work is presented in the context of smartglasses but does not assume a specific array geometry, and the framework extends to arbitrary microphone arrays.
2 Signal Model
Consider a microphone array with $N_{m}$ microphones used to capture an acoustic scene. We assume that the recorded sound field can be expressed as a superposition of signals arriving from $N_{s}$ distinct directions. In the short-time Fourier transform (STFT) domain, the signal observed at the array, at time index $t$ and frequency index $f$, is written as
$\mathbf{x}[t,f]=\mathbf{A}[f]\,\mathbf{s}[t,f]+\mathbf{v}[t,f],$ (1)
where $\mathbf{x}[t,f]\in\mathbb{C}^{N_{m}}$ is the vector of microphone signals, $\mathbf{A}[f]\in\mathbb{C}^{N_{m}\times N_{s}}$ is the acoustic transfer matrix, $\mathbf{s}[t,f]\in\mathbb{C}^{N_{s}}$ is the vector of source signals, and $\mathbf{v}[t,f]\in\mathbb{C}^{N_{m}}$ represents additive and/or diffuse noise. The dependence on time and frequency is omitted for brevity in the rest of the paper.
The second-order statistics of the signals are described by the covariance matrices
$\begin{split}\mathbf{R}_{x}&=\mathbb{E}\{\mathbf{x}\mathbf{x}^{H}\}\in\mathbb{C}^{N_{m}\times N_{m}},\qquad\\ \mathbf{R}_{s}&=\mathbb{E}\{\mathbf{s}\mathbf{s}^{H}\}\in\mathbb{C}^{N_{s}\times N_{s}}.\end{split}$ (2)
In the next section, we review both signal-dependent and independent variants of binaural signal matching. At high frequencies, the magnitude-LS variant replaces the complex LS objective while using the same $\hat{\mathbf{R}}_{s}$ construction.
3 Binaural Signal Matching
3.1 Signal-Independent Binaural Signal Matching
Signal-independent BSM aims to design a linear filter that maps the microphone array signals to binaural signals at the user’s ears. The design does not depend on a specific source signal but instead assumes a diffuse sound field. This corresponds to energy being uniformly distributed across all directions of arrival. The least-squares optimal binaural filter weights $\mathbf{c}^{\text{BSM}}\in\mathbb{C}^{2\times N_{m}}$ solve,
$\min_{\mathbf{c}}\;\left\|\mathbf{c}\mathbf{A}-\mathbf{H}^{T}\right\|_{F}^{2}$ (3)
where $\mathbf{H}\in\mathbb{C}^{N_{s}\times 2}$ denotes the head-related transfer functions (HRTFs) for the left and right ears. Under the diffuse assumption, the least-squares optimal signal-independent BSM solution takes the form
$\mathbf{c}^{\text{BSM}}=\,\mathbf{H}^{T}\,\mathbf{A}^{H}\big{(}\epsilon\mathbf{I}+\mathbf{A}\mathbf{A}^{H}\big{)}^{-1}.$ (4)
Here, $\epsilon$ is a small regularization parameter to ensure numerical stability.
3.2 Signal-Dependent Binaural Signal Matching
In this section, we introduce signal-dependent approaches to binaural signal matching. In particular, we introduce the proposed mixture of experts approach, that may be viewed as an extension of the two other methods.
3.2.1 COMPASS-BSM (COM)
COMPASS is a parametric method used to render spatial audio scenes [4]. The directions of arrival of sources are estimated first. Given those directions, the received soundfield is decomposed into direct and residual parts, where the direct part is obtained as the linearly constrained minimum variance [9] (LCMV) estimate
$\hat{\mathbf{s}}_{d}=\mathbf{W}_{d}\,\mathbf{X},\qquad\mathbf{W}_{d}=\big{(}\mathbf{A}_{d}^{H}\mathbf{R}_{x}^{-1}\mathbf{A}_{d}\big{)}^{-1}\mathbf{A}_{d}^{H}\mathbf{R}_{x}^{-1},$ (5)
where $\mathbf{A}_{d}$ stacks steering vectors of the direct component. The binaural estimate of the direct component is rendered directly through the HRTF for each corresponding source direction stacked in $\mathbf{H}_{d}$ for the two ears,
$\hat{\mathbf{p}}^{\,d}=\mathbf{H}_{d}^{T}\hat{\mathbf{s}}_{d}.$ (6)
The direct component is backprojected and subtracted from the received signal to obtain the residual
$\hat{\mathbf{X}}_{r}=\big{(}\mathbf{I}-\mathbf{A}_{d}\mathbf{W}_{d}\big{)}\mathbf{X}.$ (7)
This residual component is rendered via standard BSM weights $\mathbf{c}^{\text{BSM}}$. Therefore, the binaural residual estimate is,
$\hat{\mathbf{p}}^{\,r}=\big{(}\mathbf{c}^{\text{BSM}}\big{)}\hat{\mathbf{X}}_{r}.$ (8)
The final binaural output $\hat{\mathbf{p}}$ is obtained by summing direct and residual components and can be described via an equivalent single-stage filter:
$\begin{split}\hat{\mathbf{p}}&=\hat{\mathbf{p}}^{\,d}+\hat{\mathbf{p}}^{\,r}=\big{(}\mathbf{c}^{\text{COM}}\big{)}\mathbf{X},\qquad\\ \mathbf{c}^{\text{COM}}&=\mathbf{c}^{\text{BSM}}\big{(}\mathbf{I}-\mathbf{A}_{d}\mathbf{W}_{d}\big{)}+\mathbf{H}_{d}^{T}\mathbf{W}_{d}.\end{split}$ (9)
3.2.2 Directional BSM (d-BSM)
Form a source covariance that embeds the estimated direct component and a diffuse reverberant component[10]:
$\hat{\mathbf{R}}_{s}=\hat{\mathbf{R}}_{s_{d}}+\hat{\sigma}_{r}^{2}\mathbf{I},\qquad\hat{\mathbf{R}}_{s_{d}}=\mathbb{E}\{\hat{\mathbf{s}}_{d}\hat{\mathbf{s}}_{d}^{H}\},$ (10)
where $\hat{\sigma}_{r}^{2}$ is obtained from the residual consistent with the diffuse assumption in the paper. Let $\mathbf{A}=[\,\mathbf{A}_{d}\ \mathbf{A}_{r}\,]$ stack the direct and reverberant steering matrices and let $\mathbf{h}$ denote the ear-specific HRTF vector. The signal-dependent BSM filter is
$\mathbf{c}^{\text{d-BSM}}=\mathbf{H}^{T}\,\ \hat{\mathbf{R}}_{s}\,\mathbf{A}^{H}\big{(}\mathbf{A}\hat{\mathbf{R}}_{s}\mathbf{A}^{H}+\mathbf{R}_{n}\big{)}^{-1},$ (11)
with $\mathbf{R}_{n}$ the noise covariance.
3.3 Proposed Method
We view the binaural rendering problem as a mixture of experts [11, 12], where each expert corresponds to a binaural filter designed under the assumption of a single source from candidate directions. Let the set of candidate directions be indexed by $q=1,\ldots,Q$. For each direction $q$, we compute a signal dependent binaural filter $\mathbf{c}_{q}[f]$, assuming the sound source originates from that direction, and apply it to the microphone signals $\mathbf{X}[t,f]$ to obtain
$\hat{\mathbf{p}}_{q}[t,f]=\mathbf{c}_{q}^{H}[f]\,\mathbf{X}[t,f].$ (12)
The final output is a convex combination of the experts,
$\hat{\mathbf{p}}[t,f]=\sum_{q=1}^{Q}\alpha_{q}[t,f]\,\hat{\mathbf{p}}_{q}[t,f],$ (13)
with nonnegative weights satisfying $\sum_{q}\alpha_{q}[t,f]=1$ for each frequency $f$.
Following the online convex optimization framework, we update the weights with exponential weighting. Let $L_{q}[t,f]$ represent the cumulative loss of expert $q$ at time $t$ and frequency $f$. The blend weights are then given by
$\alpha_{q}[t,f]=\frac{\exp\!\left(-\eta L_{q}[t,f]\right)}{\sum_{j=1}^{Q}\exp\!\left(-\eta L_{j}[t,f]\right)},$ (14)
where $\eta>0$ is a learning rate. This scheme ensures that experts with lower cumulative loss contribute more strongly to the final output while preserving smooth adaptation. Additionally, it ensures that our method is time-adaptive despite using linear-time invariant filters in its construction. The instantaneous loss $\ell_{q}[t,f]$ is defined as the squared norm of the residual, summed over all microphones $N_{m}$. This loss is accumulated over time into the cumulative loss $L_{q}[t,f]$, which is updated recursively. This yields an expression for the the binaural filters produced by the proposed method.
$\mathbf{r}_{q}[t,f]$ $=\mathbf{X}[t,f]-\mathbf{A}_{q}[f]\,\hat{s}_{q}[t,f],\forall q\in{1,..,Q}$ (15)
$\ell_{q}[t,f]$ $=\sum^{N_{m}}\|\mathbf{r}_{q}[t,f]\|^{2},$ (16)
$L_{q}[t,f]$ $=L_{q}[t-1,f]+\ell_{q}[t,f],$ (17)
$\mathbf{c}^{MoE}[t,f]$ $=\sum_{q=1}^{Q}\alpha_{q}[t,f]\,\mathbf{c}_{q}[t,f],$ (18)
$\hat{\mathbf{p}}[t,f]$ $=\big{(}\mathbf{c}^{\text{MoE}}\big{)}^{H}[t,f]\,\mathbf{X}[t,f].$ (19)
Intuitively, this is like an implicit DOA estimator that relies on the assumption that the beamformer with the lowest residual likely contains the direct source.
3.3.1 Regret Bound
The exponential weighting scheme enjoys a standard regret guarantee in the online convex optimization setting. Define the regret after $T$ time frames as
$R_{T}=\sum_{t=1}^{T}\sum_{f}\ell\big{(}\hat{\mathbf{p}}[t,f]\big{)}-\min_{q}\sum_{t=1}^{T}\sum_{f}\ell\big{(}\hat{\mathbf{p}}_{q}[t,f]\big{)},$ (20)
that is, the cumulative loss of the mixture compared to the best single expert in hindsight. With $\ell(\cdot)$ convex and bounded, the regret of the exponential weighting algorithm satisfies
$R_{T}=\mathcal{O}\big{(}\sqrt{T\log Q}\big{)}.$ (21)
This guarantee implies that, asymptotically, the mixture performs nearly as well as the best expert chosen with full hindsight, while retaining the flexibility to adapt to time-varying acoustic scenes. Note that the average regret $R_{T}/T$ goes to zero as time approaches infinity. [13]
4 Field of View Enhancement
We now describe two control strategies for field of view (FoV) enhancement. Each strategy modifies the binaural signal matching (BSM) formulation to emphasize directions within a user-selected field of view while attenuating those outside it. Both signal-independent and signal-dependent variants are presented.
4.1 Gain Control
In gain control, we apply multiplicative gains to the HRTFs [8]. For each direction $q$, the modified HRTF is
$\tilde{\mathbf{h}}_{q}=\begin{cases}\mathbf{h}_{q},&q\in\text{FoV},\\ (1-\gamma)\,\mathbf{h}_{q},&q\notin\text{FoV},\end{cases}$ (22)
where $\gamma\in[0,1]$ is a user-selected gain parameter. Note, that for COMPASS we can add gain $\geq 1$ in the FoV directions that are rendered directly through the HRTF as shown in equation 6.
4.2 Distortion Control
In distortion control, we penalize deviations in the BSM matching differently for FoV and non-FoV directions. Define the diagonal distortion weighting matrix
$\mathbf{D}=\operatorname{diag}(w_{1},\ldots,w_{N_{s}}),\qquad w_{q}=\begin{cases}1,&q\in\text{FoV},\\ 1-\delta,&q\notin\text{FoV},\end{cases}$ (23)
with $\delta\in[0,1]$. The weighted least-squares BSM formulation can be written compactly for as
$\min_{\mathbf{c}}\;\big{(}\mathbf{c}\mathbf{A}-\mathbf{H}\big{)}\mathbf{D}\big{(}\mathbf{c}\mathbf{A}-\mathbf{H}\big{)}^{T},$ (24)
In this form, FoV directions are matched exactly while non-FoV directions are permitted distortion proportional to $1-\delta$.
In directional BSM (d-BSM), the distortion weights enter the covariance matrices. The weighted source covariance is
$\tilde{\mathbf{R}}_{s}=\mathbf{D}^{1/2}\mathbf{R}_{s}\mathbf{D}^{1/2},$ (25)
and the corresponding d-BSM filter becomes
$\mathbf{c}=\mathbf{H}^{T}\,\tilde{\mathbf{R}}_{s}\,\mathbf{A}^{H}\big{(}\mathbf{A}\,\tilde{\mathbf{R}}_{s}\,\mathbf{A}^{H}+\mathbf{R}_{n}\big{)}^{-1},$ (26)
4.3 Proposed FoVE via Mixture of Experts
The mixture of experts framework described earlier can be directly extended to field of view enhancement. Each expert corresponds to a binaural filter computed under a FoV-aware formulation, and the adaptive blending is performed in the same manner as before. In practice, the proposed method can be made FoV-enhanced simply by constructing the expert filters using any of the FoV-aware versions of BSM, COMPASS-BSM, or d-BSM. The resulting mixture the emphasis on user-selected FoV regions. Figure 2 depicts the direction dependent gain from varying the two parameters $\gamma$ and $\delta$ for signal independent beamforming. The same results extends to each of the signal-dependent methods relative the weights without FoVE. Note, that we elect to use magnitude least-squares optimization or MagLS. [14, 15, 2]
Fig. 2: The direction dependent gain from field of view in the frontal direction. The plot on left shows the effect of varying $\gamma$ with a fixed value of $\delta$ while the second plot shows the effect of varying $\delta$ with fixed value of $\gamma$
5 Results
5.1 Simulation
A continuous motion simulation is created in pyroomacoustics [16] in a room of size [8m, 8m, 5m] with RT60 $\approx$ 200ms. A 4-microphone array is simulated with the reference center at [4m, 4m, 2m]. All signals are sampled synchronously at 48 kHz and the talker speech is derived from the EARS dataset [17]. One moving talker is simulated with a 6-degree spacing in azimuthal direction. The talker is initialized to position [7m, 4m, 2m] in the frontal direction of the microphone array. The talker covers every 6 degrees in 167ms, which equates to a speed $\approx$ 2 $m/s$. Figure 3 demonstrates the efficacy of the proposed residual based loss, that is clearly able to track the talker in the volume. Additionally, we have depicted the value of blend weights for the set of BSM filters, that shows that in continuous motion the algorithm prefers to blend different filters resulting in perceptually relevant directions being rendered correctly resulting in accurate interaural cues at each time.
Fig. 3: The argmin of the residual energy at each time and the corresponding blend weights for the grid of beamformers for the simulation
5.2 Experimental
The simulation experiment is recreated in a real-world environment with a 4 microphone headworn microphone array sampled at 48 kHz. One talker was asked to start in front of the listener and walked in a counter-clockwise direction similar to the simulation experiment. Figure 4 demonstrates the effectiveness of the proposed method in tracking the talker as they walk in the environment. Depicted in the same figure is the blend weights of the different candidate binauralization filters over time. This clearly demonstrates that we are able to render the source with the correct binaural cues as they move through the environment especially with MoE module that would render directly using the HRTFs.
Fig. 4: The argmin of the residual energy at each time and the corresponding blend weights for the grid of beamformers in the experiment performed
6 Conclusion
In this work, a novel mixture of experts framework is theorized for binauralization. The proposed framework extends previous work in signal-dependent binauralization to scenarios with continuous motion and for adjustable field of view enhancement. Our results demonstrate that the framework is not only effective but highly modular, giving a user fidelity over the reproduced acoustic scene. Our approach leverages a softmax function to seamlessly blend outputs from numerous BSM filters, each tailored to distinct direction. This modularity enables real-time adaptation to listener movement and environmental changes, ensuring consistent spatial audio quality. Future work will focus on neural beamspace projection.